OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 7x + 8y = 51 7x + 10y = 55 Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 7x + 8y = 51 7x + 10y = 55 @Mathematics

OpenStudy (anonymous):

solve the first equation for x or y. 7x = 51-8y x = (51-8y)/7 plug x into second equation and solve for y then use y to solve for x, this will get you started

OpenStudy (anonymous):

That is quite a convoluted way to think about it. Simply subtract the two equations. You get 7x-7x+8y-10y=51-55. You can then solve for y and then by plugging away its value into one of the original equations, solve for y. Good luck.